Session: Feb 2014 (session 1)
Wavelength: 18 cm
To give feedback on this experiment, click
(Geert Kuper at Sat Feb 22 09:13:44 2014) Success:
(Orlati at Mon Feb 24 08:26:35 2014) Success:
Antabs and log files ready on EVN repository
(K. Borkowski at Mon Feb 24 10:14:15 2014) Minor failures: LATE
15 minute late start. No other known problems.
(Jun Yang at Mon Mar 03 11:21:54 2014) Success:
It was a cloudy day, while the experiment was successfully done in the e-EVN mode.